OpenStudy (anonymous):

what is the value of x? hypotenuse=26 short leg=13 (right triangle)

OpenStudy (zzr0ck3r):

 a^2 +13^2 = 26^2

OpenStudy (anonymous):

so whats my answer?

OpenStudy (zzr0ck3r):

solve for a, we are not here to give answers. we are here to help:)

OpenStudy (zzr0ck3r):

a^2=26^2-13^2

OpenStudy (anonymous):

what does those arrows mean

OpenStudy (zzr0ck3r):

\[a^2\] = a^2

OpenStudy (zzr0ck3r):

so you have to solve for a \[a^2 = 26^2-13^2\]

OpenStudy (anonymous):

my answer should be the square root of 507 right

OpenStudy (zzr0ck3r):

yes, and you would take the positive square root because the negative one does not make sense in our context

OpenStudy (anonymous):

thanks so much .. first time trying this, will be coming back all the time!

OpenStudy (zzr0ck3r):

good deal:)

OpenStudy (anonymous):

The area of a triangle inscribed in a circle is 60 cm^2 and the radius of the circumscribed circle is 8 cm. If the two sides of the inscribed triangle are 10 cm and 12 cm, respectively, find the third side. please help